# Decentralized Zeno-Free Event-Triggered Control For Multiple Networks Subject to Stochastic Network Delays and Poisson Pulsing Attacks

### v1

@@ III-A Constraints
Note that items 1)-3) do not make any reference to the NCS (II-C)-(4), and they capture the intrinsic properties of the access protocol itself, meaning that the discrete-time system (7) or the protocol is uniformly globally exponentially stable when we refer to the reset system (7) as a discrete-time system induced by the protocol $h_{i}$, see [13, Definition 7]. For the reset system (7), item 2) characterizes the reset at a transmission instant ($l_{i}$ resets from 0 to 1 and $k_{i}$ increases 1), after which the protocol works for the transmission within the $i$th network during $(t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$; while item 3) is relaxed ($\lambda_{i}=1$), since it is for the reset at a successful/faied update instant ($l_{i}$ resets from 1 to 0 and $k_{i}$ remains unchanged), after which the $i$th network is idle during $(t^{i}_{j}+\tau^{i}_{j},t^{i}_{j+1}]$ or $(t^{i}_{j}+\tau^{i}_{j},t^{i}_{j^{\prime}}]$ (see Fig. 3). That is, the reset system (7) will undergo two different modes (on-working and idle), and this is why two cases $l_{i}=\{0,1\}$ need to be applied here.
@@ III-B Stability Guarantee
In all that follows, define $\mathcal{V}:=\cup_{\omega\in\Omega,i\in\mathbb{Z}_{\geq 1}}\textbf{v}_{i}(\omega)$. A function $\mathcal{U}:\hbox{dom}\mathcal{U}\rightarrow\mathbb{R}$ is a certification candidate for $\mathcal{H}$ (II-B), denoted as $\mathcal{U}\in\mathcal{D}(\mathcal{H})$, if
@@
Based on Assumption 1, any upper semicontinuous function for $\mathcal{H}$ that satisfies C1-C2 is a certification candidate for $\mathcal{H}$, see [1, Lem. 4.1]. Sharper than upper semicontinuity, the certification candidate $\mathcal{U}$ can be partially Lipschitz for $\mathcal{H}$ if $\mathcal{U}$ is locally Lipschitz on an open set containing $C\setminus L_{\mathcal{U}}(0)$ as well as continuous on $C$, where $L_{\mathcal{U}}(0):=\{x\in\hbox{dom}\mathcal{U}:\mathcal{U}(x)=0\}$. We note that the certification candidate considered here is locally Lipschitz with respect to the flow set $C$ (due to the existence of several key instants, e.g., $\tau^{i}_{e}=\tau^{1,i}_{miet}$), rather than continuously differentiable; as shown in Assumption 4, even if “locally Lipschitz” is assumed rather than “continuously differentiable” (see [26, Lemmas 5-6], [27, Lemmas 1-2]), we can still use the standard directional derivative for a.a. $\tilde{x}\in\mathbb{R}^{n_{\tilde{x}}}$ in Assumption 4-6), or in the later computation using Hybrid Dynkin’s formula [16, Lem. 9] since the locally Lipschitz candidate is continuously differentiable when we consider it in segments (with several key instants as segmentation points, e.g., $\tau^{i}_{e}=\tau^{1,i}_{miet}$). Next we present stability properties of the closed set $\mathcal{A}$ for the SHS (II-B) with (10) and (11).
@@
Proposition 2. Let Assumptions 1-4 and Conditions 1-4 hold for the SHS (II-B) with (10) and (11). Then there exists a certification candidate $\mathcal{U}(x)$ and $\underline{\beta}_{\mathcal{U}},\overline{\beta}_{\mathcal{U}}\in\mathcal{K}_{\infty}$, such that $\underline{\beta}_{\mathcal{U}}(|x|_{\mathcal{A}})\leq\mathcal{U}(x)$ holds for $x\in C\cup D\cup G(D\times\mathcal{V})$, and $\mathcal{U}(x)\leq\overline{\beta}_{\mathcal{U}}(|x|_{\mathcal{A}})$ holds for $x\in C\cup D$.
@@
Proof: Now we show that there exists a certification candidate which is also partially Lipschitz, relative to the closed set $\mathcal{A}$. For $x\in C\cup D$, define $\mathcal{U}(x):=V(\tilde{x})+\sum^{N}_{i=1}(\gamma_{l_{i},i}\phi_{l_{i},i}W^{2}_{i}+\chi_{i})$, then $\mathcal{U}$ is a certification candidate, i.e., $\mathcal{U}(x)\in\mathcal{D}(\mathcal{H})$. Indeed, it holds from (11) that $G(D\times\mathcal{V})\subset\mathbb{X}$, which implies C1 holds, i.e., $C\cup D\cup G(D\times\mathcal{V})\subset\hbox{dom}\mathcal{U}$; as well, C2 holds from the fact that $\mathcal{U}(x)=0$ iff $x\in\mathcal{A}$, i.e., $L_{\mathcal{U}}(0)=\mathcal{A}$, and $\mathcal{U}(\mathbb{X}\setminus\mathcal{A})>0$. Assumption 4 and Condition 4 imply that function $\mathcal{U}(x)$ is locally Lipschitz on an open set containing $C\setminus L_{\mathcal{U}}(0)$ and continuous on its domain; based on Assumption 1, [1, Lem. 4.1] implies that any upper semicontinuous (weaker than locally Lipschitz) function for $\mathcal{H}$ that satisfies C1-C2 is a certification candidate for $\mathcal{H}$. Indeed, in C3 holds due to the upper semicontinuous function $\mathcal{U}(\cdot)$: $-\mathcal{U}(\cdot)$ is a normal integrand since it is upper semicontinuous [14, Exam. 14.30], and in terms of [14, Exam. 14.32&Thm. 14.37], the measurability of $v\mapsto G(x,v)$ and the outer semicontinuity of $x\mapsto G(x,v)$ in Assumption 1 (Fig. 1) imply that the quantity $\int_{\mathbb{R}^{m}}\sup_{g\in G(x,v)}\mathcal{U}(g)\mu(dv)$ is well-defined for each $x\in D$. Therefore, the partially Lipschitz function $\mathcal{U}(x)$ is a certification candidate relative to the closed set $\mathcal{A}$ for $\mathcal{H}$. Then we show the radial unboundedness of $\mathcal{U}(x)$. Define $\underline{\beta}_{\mathcal{U}}(|x|):=\min\{\underline{\alpha}(\frac{|\tilde{x}|}{2}),\sum^{N}_{i=1}\gamma_{l_{i},i}\lambda_{i}\underline{c}^{2}_{i}(\frac{|(e_{i},s_{i})|}{2})\}+\sum^{N}_{i=1}\ln(1+\chi_{i})$ and $\overline{\beta}_{\mathcal{U}}(|x|):=\max\{2\overline{\alpha}(|\tilde{x}|),\sum^{N}_{i=1}2\frac{\gamma_{l_{i},i}}{\lambda_{i}}\bar{c}^{2}_{i}(|(e_{i},s_{i})|)\}+\sum^{N}_{i=1}(\exp(\chi_{i})-1)$. Then it holds in vies of Assumption 4 for all $x\in\mathbb{X}$ that $\underline{\beta}_{\mathcal{U}}(|x(t,j)|_{\mathcal{A}})\leq\mathcal{U}(x)$ for $x\in C\cup D\cup G(D\times\mathcal{V})$, while $\mathcal{U}(x)\leq\overline{\beta}_{\mathcal{U}}(|x(t,j)|_{\mathcal{A}})$ for $x\in C\cup D$. In view of Assumption 4 and $\ln(\cdot),\exp(\cdot)\in\mathcal{K}_{\infty}$, functions $\underline{\beta}_{\mathcal{U}}$ and $\overline{\beta}_{\mathcal{U}}$ are of class $\mathcal{K}_{\infty}$. As a result, $\mathcal{U}(x)$ is a partially Lipschitz and radially unbounded certification candidate relative to the closed set $\mathcal{A}$ for the SHS (II-B) with (10) and (11).∎
@@
Theorem 1. For the SHS (II-B) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally stable in probability.
@@
with $\hat{\beta}:=\inf_{i\in\bar{N}}\{\hat{\beta}_{i}\}$. In view of [16, (10)] and Proposition 2, it holds from (29) that $\mathbb{P}(\mbox{graph}(\textbf{x})\subseteq\mathbb{R}^{2}\times(\mathcal{A}+\varepsilon\mathbb{B}))\geq 1-\overline{\beta}_{\mathcal{U}}(\delta)/\underline{\beta}_{\mathcal{U}}(\varepsilon)$. Define $\overline{\beta}_{\mathcal{U}}(\delta)/\underline{\beta}_{\mathcal{U}}(\varepsilon):=\rho$, then it holds for $\forall$ $\varepsilon,\rho>0$, $\exists~{}\delta>0$, as well as for $\forall$ $\delta,\rho>0$, $\exists~{}\varepsilon>0$ that $\textbf{x}\in\mathcal{S}_{r}(\mathcal{A}+\delta\mathbb{B})\Longrightarrow$ $\mathbb{P}(\mbox{graph}(\textbf{x})\subseteq\mathbb{R}^{2}\times(\mathcal{A}+\varepsilon\mathbb{B}))\geq 1-\rho$. Consequently, Definition 1-(a) implies the UGSp of closed set $\mathcal{A}$ for the SHS (II-B) with (10) and (11).
Theorem 2. For the SHS (II-B) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally attractive in probability.
@@
Theorem 3. For the SHS (II-B) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally asymptotically stable in probability.
@@ IV Simulations
Let $J_{i}:=\mbox{diag}\{0.13,0.13,0.04\}$, $K_{i}:=0.013I$, $\kappa_{i}:=3$ and $\underline{\delta}_{i}:=0.45$. Let $q_{d}(0):=(1,0,0,0)$ and $\Theta_{d}(0):=0$ with $\dot{\Theta}_{d}\equiv 0$. Define initial values as $q_{1}(0):=(-0.2,\sqrt{1-0.04}\vec{\upsilon})$, $q_{2}(0):=(0.2,-\sqrt{1-0.04}\vec{\upsilon})$, $q_{3}(0):=(0.2,\sqrt{1-0.04}\vec{\upsilon})$, $q_{4}(0):=(-0.2,-\sqrt{1-0.04}\vec{\upsilon})$, and $\Theta_{i}(0):=0.2\vec{\upsilon}$, where $\vec{\upsilon}:=(1,2,3)/\sqrt{14}$. In view of [13], we apply the standard sampled-data protocol with $h_{i}=0$ for $i\in\{1,2,3,4\}$, considering the strong nonlinearity of attitude kinematic and dynamic systems. Define $e_{i}:=(\hat{\tilde{q}}_{i}-\tilde{q}_{i},\hat{\tilde{\Theta}}_{i}-\tilde{\Theta}_{i})$, $\Pi_{i}:=\left(\begin{array}[]{c}-\tilde{\epsilon}^{T}_{i}\\ \tilde{\eta}_{i}I+S(\tilde{\epsilon}_{i})\end{array}\right)$, and $\tilde{x}:=(\tilde{q}_{i},J_{i}\tilde{\Theta}_{i},h_{i})_{i\in\bar{N}}\in\mathbb{X}^{\prime}$ with $\mathbb{X}^{\prime}:=(\mathcal{S}^{3})^{N}\times(\mathbb{R}^{3})^{N}\times\{-1,1\}^{N}$. Let $e_{i,1}$ denote the $\tilde{q}_{i}$-component of $e_{i}$, $e_{i,11}$ denote the $\tilde{\eta}_{i}$-component of $e_{i,1}$, $e_{i,12}$ denote the $\tilde{\epsilon}_{i}$-component of $e_{i,1}$, and $e_{i,2}$ denote the $\tilde{\Theta}_{i}$-component of $e_{i}$, i.e., $e_{i}=((e_{i,11},e_{i,12}),e_{i,2})$. The goal is to stabilize $\mathcal{A}:=(\mathcal{A}_{p,i})_{i\in\bar{N}}\times\{e=\textbf{0}_{N}\}$, where $\mathcal{A}_{p,i}:=\{\tilde{x}\in\mathbb{X}^{\prime}|\tilde{\eta}_{i}=\overline{\mbox{sgn}}(h_{i}),\tilde{\Theta}_{i}=0\}$.
@@
Choose non-negative definite functions $W_{i}:=|e_{i}|$ and $V(\tilde{x}):=\sum^{N}_{i=1}(2\kappa_{i}(1-h_{i}\tilde{\eta}_{i})+\frac{1}{2}\tilde{\Theta}^{T}_{i}J_{i}\tilde{\Theta}_{i})$. In view of Proposition 2, Assumptions 4-1) and 4-5) hold by choosing concave function $\ln(\cdot)\in\mathcal{K}_{\infty}$ and convex function $\exp(\cdot)\in\mathcal{K}_{\infty}$. For $W_{i}=|e_{i}|$ which satisfies $W_{i}=\sqrt{e^{T}_{i}e_{i}}$, it holds
$\begin{split}&\frac{\partial W_{i}}{\partial e_{i}}\cdot\dot{e}_{i}=\frac{1}{2}\frac{2e^{T}_{i}\dot{e}_{i}}{\sqrt{e^{T}_{i}e_{i}}}=\frac{e^{T}_{i}\dot{e}_{i}}{W_{i}}=\frac{e^{T}_{i}}{W_{i}}g(\tilde{x},e_{i})\\ =&\frac{e^{T}_{i}}{W_{i}}[\frac{1}{2}\Pi_{i}*(-h_{i}\hat{\tilde{\epsilon}}_{i}),-J^{-1}_{i}(\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})\tilde{\Theta}_{i}\\ &-\kappa_{i}h_{i}\hat{\tilde{\epsilon}}_{i}-K_{i}\hat{\tilde{\Theta}}_{i})]\\ =&\frac{e^{T}_{i}}{W_{i}}[\frac{1}{2}\Pi_{i}(-h_{i}(\tilde{\epsilon}_{i}+e_{i,12})),-J^{-1}_{i}(\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})\tilde{\Theta}_{i}\\ &-\kappa_{i}h_{i}(\tilde{\epsilon}_{i}+e_{i,12})-K_{i}(\tilde{\Theta}_{i}+e_{i,2}))]\\ =&\frac{e_{i,1}^{T}}{W_{i}}(-\frac{1}{2}\Pi_{i}h_{i}\tilde{\epsilon}_{i}-\frac{1}{2}\Pi_{i}h_{i}e_{i,12})\\ &+\frac{e_{i,2}^{T}}{W_{i}}[-J^{-1}_{i}(\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})\tilde{\Theta}_{i}\\ &-\kappa_{i}h_{i}\tilde{\epsilon}_{i}-\kappa_{i}h_{i}e_{i,12}-K_{i}\tilde{\Theta}_{i}-K_{i}e_{i,2})]\\ \leq&\frac{1}{2}|h_{i}\tilde{\epsilon}_{i}|+\frac{1}{2}|h_{i}e_{i,12}|+|J^{-1}_{i}||\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})||\tilde{\Theta}_{i}|\\ &+J^{-1}_{i}\kappa_{i}|\tilde{\epsilon}_{i}|+J^{-1}_{i}K_{i}|\tilde{\Theta}_{i}|+|\kappa_{i}J^{-1}_{i}e_{i,12}|+J^{-1}_{i}K_{i}|e_{i,2}|\\ \leq&\frac{1}{2}|h_{i}e_{i,12}|+|\kappa_{i}J^{-1}_{i}e_{i,12}|+|J^{-1}_{i}K_{i}e_{i,2}|\\ &+\frac{1}{2}|h_{i}\tilde{\epsilon}_{i}|+J^{-1}_{i}\kappa_{i}|\tilde{\epsilon}_{i}|+|J^{-1}_{i}K_{i}\tilde{\Theta}_{i}|\\ \leq&L_{l_{i},i}W_{i}+H_{l_{l},i},\end{split}$
where $|\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})|=0$, $L_{l_{i},i}\geq\max\{\frac{1}{2}+J^{-1}_{i}\kappa_{i},|J^{-1}_{i}K_{i}|\}=\max\{0.65,0.2\}=0.65$ and $H_{l_{i},i}:=(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})|\tilde{\epsilon}_{i}|+|J^{-1}_{i}K_{i}\tilde{\Theta}_{i}|=0.65|\tilde{\epsilon}_{i}|+0.2|\tilde{\Theta}_{i}|$, then Assumption 4-4) holds. For $V(\tilde{x})=\sum^{N}_{i=1}(2\kappa_{i}(1-h_{i}\tilde{\eta}_{i})+\frac{1}{2}\tilde{\Theta}^{T}_{i}J_{i}\tilde{\Theta}_{i})$, it holds
@@
$\begin{split}&\dot{V}(\tilde{x})\\ =&\sum^{N}_{i=1}(-2\kappa_{i}h_{i}\dot{\tilde{\eta}}_{i}+\tilde{\Theta}^{T}_{i}(\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})\tilde{\Theta}_{i}-\kappa_{i}h_{i}\hat{\tilde{\epsilon}}_{i}-K_{i}\hat{\tilde{\Theta}}_{i}))\\ =&\sum^{N}_{i=1}(-\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}(\tilde{\epsilon}_{i}+e_{i,12})+0\\ &-\kappa_{i}h_{i}\tilde{\Theta}^{T}_{i}(\tilde{\epsilon}_{i}+e_{i,12})-\tilde{\Theta}^{T}_{i}K_{i}(\tilde{\Theta}_{i}+e_{i,2}))\\ =&\sum^{N}_{i=1}(-\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}-\kappa_{i}h_{i}\tilde{\Theta}^{T}_{i}\tilde{\epsilon}_{i}-\tilde{\Theta}^{T}_{i}K_{i}\tilde{\Theta}_{i}\\ &-\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}e_{i,12}-\kappa_{i}h_{i}\tilde{\Theta}^{T}_{i}e_{i,12}-\tilde{\Theta}^{T}_{i}K_{i}e_{i,2})\\ =&\sum^{N}_{i=1}(-\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}-\tilde{\Theta}^{T}_{i}K_{i}\tilde{\Theta}_{i}+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ &+\kappa_{i}((h_{i})^{2}|\tilde{\epsilon}_{i}|+|h_{i}|\tilde{\Theta}^{2}_{i})|e_{i,12}|+\lambda_{\max}(K_{i})|\tilde{\Theta}_{i}||e_{i,2}|)\\ =&\sum^{N}_{i=1}(-\frac{1}{10}(\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\tilde{\Theta}^{T}_{i}K_{i}\tilde{\Theta}_{i})\\ &-\frac{9}{10}(\kappa_{i}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\lambda_{\min}(K_{i})\tilde{\Theta}^{T}_{i}\tilde{\Theta}_{i})+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ &+\kappa_{i}((h_{i})^{2}|\tilde{\epsilon}_{i}|+|h_{i}|\tilde{\Theta}^{2}_{i})|e_{i,12}|+\lambda_{\max}(K_{i})|\tilde{\Theta}_{i}||e_{i,2}|)\\ \leq&-\tilde{\rho}V(\tilde{x})+\sum^{N}_{i=1}(-\tilde{\varrho}_{i}(x_{p,i})-H_{l_{i},i}^{2}-\sigma_{l_{i},i}(W_{i})+\gamma^{2}_{l_{i},i}W^{2}_{i}),\end{split}$

### v3

@@ III-A Constraints
Note that items 1)-3) do not make any reference to the NCS (2)-(4), and they capture the intrinsic properties of the access protocol itself, meaning that the discrete-time system (7) or the protocol is uniformly globally exponentially stable when we refer to the reset system (7) as a discrete-time system induced by the protocol $h_{i}$, see [13, Definition 7]. For the reset system (7), item 2) characterizes the reset at a transmission instant ($l_{i}$ resets from 0 to 1 and $k_{i}$ increases 1), after which the protocol works for the transmission within the $i$th network during $(t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$; while item 3) is relaxed ($\lambda_{i}=1$), since it is for the reset at a successful/faied update instant ($l_{i}$ resets from 1 to 0 and $k_{i}$ remains unchanged), after which the $i$th network is idle during $(t^{i}_{j}+\tau^{i}_{j},t^{i}_{j+1}]$ or $(t^{i}_{j}+\tau^{i}_{j},t^{i}_{j^{\prime}}]$ (see Fig. 3). That is, the reset system (7) will undergo two different modes (on-working and idle), and this is why two cases $l_{i}=\{0,1\}$ need to be applied here.
@@ III-B Stability Guarantee
In all that follows, define $\mathcal{V}:=\cup_{\omega\in\Omega,i\in\mathbb{Z}_{\geq 1}}\textbf{v}_{i}(\omega)$. A function $\mathcal{U}:\hbox{dom}\mathcal{U}\rightarrow\mathbb{R}$ is a certification candidate for $\mathcal{H}$ (1), denoted as $\mathcal{U}\in\mathcal{D}(\mathcal{H})$, if
@@
Based on Assumption 1, any upper semicontinuous function for $\mathcal{H}$ that satisfies C1-C2 is a certification candidate for $\mathcal{H}$, see [1, Lem. 4.1]. Sharper than upper semicontinuity, the certification candidate $\mathcal{U}$ can be partially Lipschitz for $\mathcal{H}$ if $\mathcal{U}$ is locally Lipschitz on an open set containing $C\setminus L_{\mathcal{U}}(0)$ as well as continuous on $C$, where $L_{\mathcal{U}}(0):=\{x\in\hbox{dom}\mathcal{U}:\mathcal{U}(x)=0\}$. We note that the certification candidate considered here is locally Lipschitz with respect to the flow set $C$ (due to the existence of several key instants, e.g., $\tau^{i}_{e}=\tau^{1,i}_{miet}$), rather than continuously differentiable; as shown in Assumption 4, even if “locally Lipschitz” is assumed rather than “continuously differentiable” (see [26, Lemmas 5-6], [27, Lemmas 1-2]), we can still use the standard directional derivative for a.a. $\tilde{x}\in\mathbb{R}^{n_{\tilde{x}}}$ in Assumption 4-6), or in the later computation using Hybrid Dynkin’s formula [16, Lem. 9] since the locally Lipschitz candidate is continuously differentiable when we consider it in segments (with several key instants as segmentation points, e.g., $\tau^{i}_{e}=\tau^{1,i}_{miet}$). Next we present stability properties of the closed set $\mathcal{A}$ for the SHS (1) with (10) and (11).
@@
Proposition 2. Let Assumptions 1-4 and Conditions 1-4 hold for the SHS (1) with (10) and (11). Then there exists a certification candidate $\mathcal{U}(x)$ and $\underline{\beta}_{\mathcal{U}},\overline{\beta}_{\mathcal{U}}\in\mathcal{K}_{\infty}$, such that $\underline{\beta}_{\mathcal{U}}(|x|_{\mathcal{A}})\leq\mathcal{U}(x)$ holds for $x\in C\cup D\cup G(D\times\mathcal{V})$, and $\mathcal{U}(x)\leq\overline{\beta}_{\mathcal{U}}(|x|_{\mathcal{A}})$ holds for $x\in C\cup D$.
@@
Proof: Now we show that there exists a certification candidate which is also partially Lipschitz, relative to the closed set $\mathcal{A}$. For $x\in C\cup D$, define $\mathcal{U}(x):=V(\tilde{x})+\sum^{N}_{i=1}(\gamma_{l_{i},i}\phi_{l_{i},i}W^{2}_{i}+\chi_{i})$, then $\mathcal{U}$ is a certification candidate, i.e., $\mathcal{U}(x)\in\mathcal{D}(\mathcal{H})$. Indeed, it holds from (11) that $G(D\times\mathcal{V})\subset\mathbb{X}$, which implies C1 holds, i.e., $C\cup D\cup G(D\times\mathcal{V})\subset\hbox{dom}\mathcal{U}$; as well, C2 holds from the fact that $\mathcal{U}(x)=0$ iff $x\in\mathcal{A}$, i.e., $L_{\mathcal{U}}(0)=\mathcal{A}$, and $\mathcal{U}(\mathbb{X}\setminus\mathcal{A})>0$. Assumption 4 and Condition 4 imply that function $\mathcal{U}(x)$ is locally Lipschitz on an open set containing $C\setminus L_{\mathcal{U}}(0)$ and continuous on its domain; based on Assumption 1, [1, Lem. 4.1] implies that any upper semicontinuous (weaker than locally Lipschitz) function for $\mathcal{H}$ that satisfies C1-C2 is a certification candidate for $\mathcal{H}$. Indeed, in C3 holds due to the upper semicontinuous function $\mathcal{U}(\cdot)$: $-\mathcal{U}(\cdot)$ is a normal integrand since it is upper semicontinuous [14, Exam. 14.30], and in terms of [14, Exam. 14.32&Thm. 14.37], the measurability of $v\mapsto G(x,v)$ and the outer semicontinuity of $x\mapsto G(x,v)$ in Assumption 1 (Fig. 1) imply that the quantity $\int_{\mathbb{R}^{m}}\sup_{g\in G(x,v)}\mathcal{U}(g)\mu(dv)$ is well-defined for each $x\in D$. Therefore, the partially Lipschitz function $\mathcal{U}(x)$ is a certification candidate relative to the closed set $\mathcal{A}$ for $\mathcal{H}$. Then we show the radial unboundedness of $\mathcal{U}(x)$. Define $\underline{\beta}_{\mathcal{U}}(|x|):=\min\{\underline{\alpha}(\frac{|\tilde{x}|}{2}),\sum^{N}_{i=1}\gamma_{l_{i},i}\lambda_{i}\underline{c}^{2}_{i}(\frac{|(e_{i},s_{i})|}{2})\}+\sum^{N}_{i=1}\ln(1+\chi_{i})$ and $\overline{\beta}_{\mathcal{U}}(|x|):=\max\{2\overline{\alpha}(|\tilde{x}|),\sum^{N}_{i=1}2\frac{\gamma_{l_{i},i}}{\lambda_{i}}\bar{c}^{2}_{i}(|(e_{i},s_{i})|)\}+\sum^{N}_{i=1}(\exp(\chi_{i})-1)$. Then it holds in vies of Assumption 4 for all $x\in\mathbb{X}$ that $\underline{\beta}_{\mathcal{U}}(|x(t,j)|_{\mathcal{A}})\leq\mathcal{U}(x)$ for $x\in C\cup D\cup G(D\times\mathcal{V})$, while $\mathcal{U}(x)\leq\overline{\beta}_{\mathcal{U}}(|x(t,j)|_{\mathcal{A}})$ for $x\in C\cup D$. In view of Assumption 4 and $\ln(\cdot),\exp(\cdot)\in\mathcal{K}_{\infty}$, functions $\underline{\beta}_{\mathcal{U}}$ and $\overline{\beta}_{\mathcal{U}}$ are of class $\mathcal{K}_{\infty}$. As a result, $\mathcal{U}(x)$ is a partially Lipschitz and radially unbounded certification candidate relative to the closed set $\mathcal{A}$ for the SHS (1) with (10) and (11).∎
@@
Theorem 1. For the SHS (1) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally stable in probability.
@@
with $\hat{\beta}:=\inf_{i\in\bar{N}}\{\hat{\beta}_{i}\}$. In view of [16, (10)] and Proposition 2, it holds from (29) that $\mathbb{P}(\mbox{graph}(\textbf{x})\subseteq\mathbb{R}^{2}\times(\mathcal{A}+\varepsilon\mathbb{B}))\geq 1-\overline{\beta}_{\mathcal{U}}(\delta)/\underline{\beta}_{\mathcal{U}}(\varepsilon)$. Define $\overline{\beta}_{\mathcal{U}}(\delta)/\underline{\beta}_{\mathcal{U}}(\varepsilon):=\rho$, then it holds for $\forall$ $\varepsilon,\rho>0$, $\exists~{}\delta>0$, as well as for $\forall$ $\delta,\rho>0$, $\exists~{}\varepsilon>0$ that $\textbf{x}\in\mathcal{S}_{r}(\mathcal{A}+\delta\mathbb{B})\Longrightarrow$ $\mathbb{P}(\mbox{graph}(\textbf{x})\subseteq\mathbb{R}^{2}\times(\mathcal{A}+\varepsilon\mathbb{B}))\geq 1-\rho$. Consequently, Definition 1-(a) implies the UGSp of closed set $\mathcal{A}$ for the SHS (1) with (10) and (11).
Theorem 2. For the SHS (1) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally attractive in probability.
@@
Theorem 3. For the SHS (1) with (10) and (11) under Assumptions 1-4 and Conditions 1-4, the closed set $\mathcal{A}$ is uniformly globally asymptotically stable in probability.
@@ IV Simulations
Let $J_{i}:=\mbox{diag}\{0.13,0.13,0.04\}$, $K_{i}:=0.013I$, $\kappa_{i}:=3$ and $\underline{\delta}_{i}:=0.45$. Let $q_{d}(0):=(1,0,0,0)$ and $\Theta_{d}(0):=0$ with $\dot{\Theta}_{d}\equiv 0$. Define initial values as $q_{1}(0):=(-0.2,\sqrt{1-0.04}\vec{\upsilon})$, $q_{2}(0):=(0.2,-\sqrt{1-0.04}\vec{\upsilon})$, $q_{3}(0):=(0.2,\sqrt{1-0.04}\vec{\upsilon})$, $q_{4}(0):=(-0.2,-\sqrt{1-0.04}\vec{\upsilon})$, and $\Theta_{i}(0):=0.2\vec{\upsilon}$, where $\vec{\upsilon}:=(1,2,3)/\sqrt{14}$. In view of [13], we apply the standard sampled-data protocol with $h_{i}=0$ for $i\in\{1,2,3,4\}$, considering the strong nonlinearity of attitude kinematic and dynamic systems. Define $\tilde{x}:=(\tilde{q}_{i},J_{i}\tilde{\Theta}_{i},h_{i})_{i\in\bar{N}}\in\mathbb{X}^{\prime}$ with $\mathbb{X}^{\prime}:=(\mathcal{S}^{3})^{N}\times(\mathbb{R}^{3})^{N}\times\{-1,1\}^{N}$. The goal is to stabilize $\mathcal{A}:=(\mathcal{A}_{p,i})_{i\in\bar{N}}\times\{e=\textbf{0}_{N}\}$, where $\mathcal{A}_{p,i}:=\{\tilde{x}\in\mathbb{X}^{\prime}|\tilde{\eta}_{i}=\overline{\mbox{sgn}}(h_{i}),\tilde{\Theta}_{i}=0\}$.
@@
Choose non-negative definite functions $W_{i}:=|e_{i}|$ and $V(\tilde{x}):=\sum^{N}_{i=1}(2\kappa_{i}(1-h_{i}\tilde{\eta}_{i})+\frac{1}{2}\tilde{\Theta}^{T}_{i}J_{i}\tilde{\Theta}_{i})$. In view of Proposition 2, Assumptions 4-1) and 4-5) hold by choosing concave function $\ln(\cdot)\in\mathcal{K}_{\infty}$ and convex function $\exp(\cdot)\in\mathcal{K}_{\infty}$. To simplify the computations, we consider Assumption 4-6) in the absence of the network, then it holds
@@
$\begin{split}&\dot{V}(\tilde{x})\\ =&\sum^{N}_{i=1}(-2\kappa_{i}h_{i}\dot{\tilde{\eta}}_{i}+\tilde{\Theta}^{T}_{i}(\Sigma(\tilde{\Theta}_{i},\bar{\Theta}_{d,i})\tilde{\Theta}_{i}-\kappa_{i}h_{i}\tilde{\epsilon}_{i}-K_{i}\tilde{\Theta}_{i}))\\ =&\sum^{N}_{i=1}(-\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}-\tilde{\Theta}^{T}_{i}K_{i}\tilde{\Theta}_{i}+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ =&\sum^{N}_{i=1}(-\frac{1}{10}(\kappa_{i}(h_{i})^{2}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\tilde{\Theta}^{T}_{i}K_{i}\tilde{\Theta}_{i})\\ &-\frac{9}{10}(\kappa_{i}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\lambda_{\min}(K_{i})\tilde{\Theta}^{T}_{i}\tilde{\Theta}_{i})+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ \leq&-\tilde{\rho}V(\tilde{x})+\sum^{N}_{i=1}(-\tilde{\varrho}_{i}(x_{p,i})-H_{l_{i},i}^{2}-\sigma_{l_{i},i}(W_{i})+\gamma^{2}_{l_{i},i}W^{2}_{i}),\end{split}$